john's age and mary's age sum up to 48. 60years later,john's age was double that of mary.
what are thier ages?

Answer by fractalier(6550)

Call John and Mary's age NOW, j and m. Then we have
j + m = 48, or j = 48 - m
Sixty years later,
j + 60 = 2(m + 60)
Now substitute in for j and get
48 - m + 60 = 2(m + 60)
108 - m = 2m + 120
This doesn't work...I think you meant six years later...
j + 6 = 2(m + 6)
48 - m + 6 = 2(m + 6)
54 - m = 2m + 12
42 = 3m
m = 14
j = 34